What is the y-intercept of the line given by the following equation below?

11x - 7y = 121

Answer by ewatrrr(24785)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
11x - 7y = 121 0r y = (11/7) - 121/7, y-intercept is -121/7